# Fragment driven dual-stream CNN Network with attention for Offline Writer Identification using Word Data

### v2

@@ Abstract
In this paper, we propose a text-independent writer identification system utilizing a dual-stream Convolutional Neural Network (CNN) trained solely using fragments extracted from word images. More specifically, we employ the integration of feature maps from the classification and dissimilarity frameworks to extract features specific to a writer as well ass writer-independent cues, that are prevalent across fragments of several writers. This aspect makes our proposal different to previous works on convolution networks for writer identification, that are trained on writer-specific word images and their corresponding fragments. Furthermore, we explore an attention mechanism that refines the ability to discern relevant features by emphasizing important regions within the fragments of the writers. Experiments conducted on word images from the IAM, CVL, and CERUG-EN databases suggest that the proposed network architecture captures well the intricate characteristics of fragments contributed by different writers. The results obtained are promising compared to previous end-to-end deep learning frameworks for identifying writers using word images as input.
@@
Writer identification, fragment, Attention network, Convolution Neural Network.
@@ I Introduction
Handwriting is a form of behavioural biometric [1] having a distinguishing trait and has been effectively employed in fields such as forensic analysis [2], historical document analysis [3, 4], and security [5] to name a few. Writer identification involves presenting a list of likely writers in order to identify the authorship of a document from a collection of reference documents stored in the database. Based on the methodology employed to acquire the data it can be broadly grouped into offline and online [6]. In the former, the temporal attributes such as coordinates, pressure, and angle is recorded by specific devices such as a tablet. As contrast to it, in the latter, the document containing handwritten text is captured in the form of an image and analysed subsequently. Based on the content of the handwritten text, writer identification systems are categorized into text-dependent and text-independent approaches. In the text-dependent approach, the content of the handwritten text is fixed, while, in the text-independent approach, the content of the handwritten text can vary. Traditional methods for writer identification rely on extracting handcrafted features to analyze and capture the unique characteristics of a writer’s style from handwritten documents. These features are specifically designed to highlight attributes such as curvature [7], slant [8], and ink width [9], among others. Since these handcrafted features depend on utilizing statistical information from handwriting, they typically require a significant amount of textual data, such as a paragraph or a block of text containing multiple sentences [3, 10] to achieve reliable performance.
Traditional writer identification systems utilizing handcrafted features achieve high recognition rates on public datasets. However, there may be applications for which access to handwritten data is limited. In such cases, decisions may need to be made based on a very small amount of handwriting, such as individual words. This scenario presents a significant challenge, as the writer-specific style information is substantially reduced compared to page or text-level inputs. Consequently, handcrafted feature-based systems fail in these situations, yielding low recognition rates.
@@ II Research framework
Convolutional networks that rely solely on classification-based frameworks typically perform well when distinguishing between dissimilar samples. However, for fine-grained tasks such as writer identification, these approaches often struggle to capture the subtle variations between similar handwriting styles—especially when working with short word images that contain limited text. Keeping this in perspective, we introduce a dual stream approach that integrates classification and dissimilarity learning frameworks on the word fragments to achieve more robust and discriminative feature representations. While the classification branch captures localized, writer-specific features from the fragments, the dissimilarity branch focuses on learning pairwise similarities, enabling the model to distinguish fine-grained variations that lie beyond predefined class boundaries. For this, we consider utilizing a Siamese architecture that is pre-trained on a character dataset but subsequently fine-tuned using the fragment data. An integrated attention mechanism further enhances our framework by highlighting the most informative regions within each fragment. This attention-guided, learning strategy on the fragments significantly improves the ability of the model to identify subtle cues, leading to better performance in writer identification tasks.
@@
1) The proposed algorithm integrates classification and dissimilarity frameworks in a dual channel network to extract both localized, writer-specific features and writer-independent features (that go beyond predefined class boundaries).
@@ III Proposed Method
In this section, we present the details of the proposed method for writer identification. Our strategy leverages a framework based on Convolutional Neural Networks and attention mechanism. We process image fragments extracted from word images through two parallel CNN modules: one trained in a writer-dependent manner to capture writer-specific details from the fragments and the other employing a writer-independent approach to extract features of fragments that are prevalent across several writers. The combined features (post attention) are passed through a classification block with fully connected layers, which assign probability scores to the fragments using a softmax activation function. The scores are then accumulated across all fragments of a word image, and the identity is determined based on the maximum accumulated score. Further, our proposed self-attention mechanism leverages relationships within the convolutional layer feature maps, highlighting critical regions in the fragments. By generating attention weights, the network refines its ability to discern relevant features, thus improving performance in writer identification. Figure 1 provides a pictorial description of our approach.
Figure 1: Pictorial overview of our proposed system.
@@ III-A Proposed Dual-stream CNN Model
The writer-independent module is designed to extract generalized features from input fragments, that tend to remain consistent across different writing instances of the individuals enrolled in the system. These patterns can be incorporated into writer identification, particularly in unconstrained settings.
@@
The Siamese network is pretrained on the Omniglot dataset [18, 19] , which was chosen for its diverse range of character styles and its suitability for one-shot image recognition tasks  [^2] In order to support the claim on the diversity range of character styles of the Omniglot dataset, we compare its performance in sub-section IV-B with the EMNIST dataset.. Architecturally, each of the CNN pathways of the Siamese writer-independent network mirrors the writer-dependent model (presented in Figure 2(b)), with the addition of Global Average Pooling, Dropout, and Fully Connected layers after the third residual block.
The pre-training of the writer independent block is conducted using the TensorFlow framework, with the network optimized via the Adam optimizer [20] over 100 epochs. The learning rate is initialized at 0.001, and a weight decay factor of 0.5 is applied when the validation loss fails to improve over five consecutive epochs.
@@
$\hat{F}=f_{c}(F_{WD},F_{WI})$ (1)
Here $F_{WI}$ and $F_{WD}$ represents the writer-dependent and independent features, respectively, while $f_{c}$ denotes the combining function. The combined feature (F^)($\hat${F}) is obtained by applying the combining function $f_{c}$ to the feature maps extracted from the outputs of the third residual block of both branches. This combined feature is then passed to the classification block, which consists of a Global Average Pooling (GAP) layer, a Dropout layer, and a Fully Connected (FC) layer. This sequence is designed to regularize the model and improve its generalization capability. The detailed architecture of the classification block is illustrated in Figure 2(d). In our study, we have explored different methodologies to obtain an efficient combining function, the details of which is discussed in section IV.
The combined network in Figure 2 (a) is trained in an end-to-end framework, where the weights of the writer-dependent module are updated from scratch, while the writer-independent module leverages transfer learning. Specifically, the weights up to the second residual block of this module are kept frozen, and fine-tuning is performed on the subsequent layers to adapt the pre-trained parameters to the fragments of the input word.
@@ III-B Attention module
In this work, we propose an attention mechanism to selectively focus on specific elements within the fragments of a word, thus assigning varying levels of importance to different regions. This adaptability boosts the capability of the model to discern complex patterns and relationships, thus enabling the combined network of Figure 2 to learn more informative features.
@@
Our approach utilizes a MobileViT-inspired multi-head attention mechanism, outlined in [21]. The input feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ is reshaped into a two-dimensional matrix $\mathbf{X}_{U}\in\mathbb{R}^{N\times C}$, where $N=H\cdot W$ denote the spatial dimensions of X. An encoder block $G$ projects $\mathbf{X}_{U}$ into a $d$-dimensional embedding space.
@@
After reshaping to $\mathbb{R}^{H\times W\times C}$, this output is added to the original input feature map $X$, to fuse the local and global contextual representations. Figure 4 shows the pictorial overview of the attention module.
@@ III-C Implementation details
The fragments extracted from the word image serve as input to the unified model, which integrates both the writer-dependent and writer-independent modules. The writer-dependent module focuses on capturing fine-grained characteristics specific to an individual writer at the fragment level. In contrast, the writer-independent module aims to learn generalized representations of fragments by leveraging the shared characteristics contributed by all writers enrolled in the system.
During training, gradient backpropagation is guided by the label smoothing cross-entropy loss function [22], which converts the one-hot encoded hard labels $y_{i}$ into soft labels $\hat{y}_{i}$ as follows:
@@
Each word is divided into fragments of equal size. With regards to each of the fragments, they are resized to $105\times 105$ while maintaining the aspect ratio and padding (if necessary) with white pixels. These modified image patches are then passed through the dual-stream CNN module to obtain scores that are then accumulated to get the identity of the writer. In our implementation, we use nine fragments .
@@ IV-A Dataset Description
In the following subsections, we present an ablation study to evaluate the effectiveness of each module integrated into the proposed system. The contribution of individual components is analyzed through a series of experiments, highlighting their impact on the overall performance of the writer identification system.
IV-B Performance evaluation of the writer-dependent and writer independent network
@@
In the first experiment, we compare the individual performance of the baseline writer-dependent module with that of the writer-independent module. The writer-dependent module is trained end-to-end using fragments extracted from the words of a specific writer. For obtaining the identity of the user from this module alone, we consider the end to end version of Figure 2 (b) by augmenting the blocks of Global Average Pooling, Dropout, and Fully Connected layers after the last residual network.
@@
In contrast, the writer-independent network is trained using a Siamese architecture with a triplet loss function. In this ablation study, we consider the pre-training performance of this module on two separate datasets: EMNIST [27] and Omniglot [18]. Once trained, writer fragments are passed through one branch of the Siamese network. It may be worth reminding here that this network is the end to end version of Figure 2 (b) obtained by augmenting the blocks of Global Average Pooling, Dropout, and Fully Connected layers after the last residual network. The primary objective of using two different datasets for pre-training the writer-independent network is to assess the influence of dataset choice on its performance.
@@
Table II presents the performance of the baseline writer-dependent module, while Table III shows the results of the writer-independent module pre-trained on the EMNIST and Omniglot datasets, each evaluated across different dimensions of the embedded feature representations.
@@
Based on the results presented in Tables II and III, it can be observed that the baseline module trained directly on writer fragments outperforms the network trained on writer-independent datasets. It is primarily because the network trained directly on the writer fragments can capture writer-specific local features more effectively compared to the network trained on writer writer-independent dataset. Moreover, based on the result of Table III following observations can be made:
1) The performance of the network initially improves as the dimension of the embedding layer increases. This is because lower-dimensional embeddings lack the capacity to capture complex, writer-specific features effectively. As the embedding size increases, the network is better able to represent these intricate characteristics, leading to improved identification accuracy. However, beyond a certain point, further increasing the embedding dimension introduces redundant or noisy information that does not contribute meaningfully to the task. Instead of capturing additional discriminative features, the network begins to encode non-generalisable patterns, which ultimately degrade performance.
@@
2) The network trained on the Omniglot dataset outperforms the one trained on EMNIST. This is primarily because Omniglot contains handwriting samples from diverse range of chracter styles, promoting the learning of more diverse and generalised writer-independent features. This diversity reduces the risk of the network overfitting to dataset-specific characteristics, thereby enhancing its ability to generalise across different writers and writing styles.
@@
TABLE III: Average writer identification rate (in %) based on the word image fragment representations obtained from the penultimate layer of the Writer Independent network. The pre-training of the network is done on the samples of the EMNIST and Omniglot datasets with varying sizes of the penultimate layer. The best identification rate is marked in bold.
@@ IV-C Impact of the dual stream network without attention
Table IV demonstrates that integrating the writer-independent module with the baseline writer-dependent model leads to a notable improvement in overall system performance. This enhancement can be attributed to the complementary strengths of both modules. The writer-dependent component is adept at capturing fine-grained, writer-specific local features by training directly on fragments from the target writers. However, its generalization capability is often limited, and constrained by the amount of text available for training. However upon integration with the writer-independent module, the resulting combined model is able to learn generalized, invariant features that are robust to variations in writing style, character morphology, and script structure.
@@ IV-D Impact of the dual stream network with attention
We now investigate the impact of incorporating an attention module on the overall performance of the system, as discussed in section III-B. Explicitly, two configurations of the attention module presented in Fig 3 (a) and (b) are analysed.
Table V summarizes the results obtained from these configurations. Based on the entries, it may be inferred that integrating the attention mechanism impacts the overall performance of the system. Upon comparing the baseline with the attention configuration in Fig 3 (a), the identification precision shows notable improvements with accuracy increasing from 92.87 %, 90.06 %, 96.15 % to 93.92 %, 92.62 % and 97.71 %, for the IAM, CVL, and CERUG-EN datasets, respectively. Likewise, the configuration in Fig 3 (b) results in best accuracy of 94.43 %, 93.15 %, 97.71 %. The results pertain to a multi-head attention setup with 2 attention heads and a dimensionality of 64 for the query, key, and value vectors. This configuration is chosen based on observations that increasing the number of heads does not improve overall system performance. Additionally, increasing the dimensionality of the query vectors tends to degrade performance, suggesting that a compact representation is more effective in this context.
@@
IV-E Efficacy of fragment driven training of the dual network
Moving further, we would like to investigate how the baseline CNN network would perform when trained on whole word images in place of its constituent fragments. It is interesting to see that the fragment level training of the network outperforms over that of whole word images (refer Table VI). This trend may be attributed to the fact that convolution networks trained on word images make decisions based on the holistic features, which may not fully capture all the relevant details of the writing style. Contrast to that, our fragment-based training approach treats each image as a collection of local patches, thereby possibly allowing the network to learn more relevant features. This results in a robust representation and in turn explains the effectiveness of the proposed identification process.
TABLE VI: Performance of writer identification system with convolution network trained on word and fragments.
@@ IV-E Efficacy of fragment driven training of the dual network
For completeness, we also present the influence of CNNs trained on both word images and image fragments using heat map visualizations, corresponding to a sample from the IAM dataset (Figure 6). These heat maps highlight the regions of the input emphasized by the network. Specifically, column 2 shows the activation map from the model trained on complete word images, while columns (3–5) display activation maps generated by models trained on image fragments under the baseline, combined, and attention configurations, respectively. For direct comparison, the fragment-based heat maps are combined together to generate the heat covering the whole input word sample. The model trained on complete word images (column 2) produces broad activations distributed across the word, reflecting a global feature extraction strategy. In contrast, the baseline fragment-trained network (column 3)exhibits scattered activations, which is scattered aroud fragments capturing individual fragment-based stroke details. The combined configuration (column 4) achieves more continuous activations spanning both character strokes and inter-character transitions, balancing local and global information. Finally, the attention-based configuration (column 5)yields the most focused activations, closely aligning with handwriting strokes and highlighting discriminative regions, thereby demonstrating superior ability to capture writer-specific features.
Figure 6: Visualization of heatmap activations for a writer sample across different network configurations. Rows 1–3 correspond to the IAM, CVL, and CERUG-EN datasets, respectively. Column 2 shows the heatmap produced by the network trained on complete word images, while columns 3–5 display heatmaps obtained from networks trained on image fragments under the baseline, combined, and attention-based configurations.

### v3

@@ Abstract
In this paper, we propose a text-independent writer identification system utilizing a dual-stream Convolutional Neural Network (CNN) trained solely using fragments extracted from word images. More specifically, we employ the integration of feature maps from the classification and dissimilarity frameworks to extract features specific to a writer as well as cues, that are prevalent across fragments of several writers. This aspect makes our proposal different to previous works on convolution networks, that are trained on writer-specific word images and their corresponding fragments. Furthermore, we explore an attention mechanism that refines the ability to discern relevant features by emphasizing important regions within the fragments. Experiments conducted on word images from the IAM, CVL, and CERUG-EN databases suggest that the proposed network architecture captures well the intricate characteristics of fragments contributed by different writers. The results obtained are promising to previous end-to-end deep learning frameworks proposed in the literature.
@@
Writer identification, fragment, Attention module, Convolution Neural Network, Writer dependent module, Writer independent module, Transfer learning.
@@ I Introduction
Handwriting is a form of behavioural biometric [1] having a distinguishing trait and has been effectively employed in fields such as forensic analysis [2], historical document analysis [3, 4], and security [5] to name a few. Writer identification involves presenting a list of likely writers in order to identify the authorship of a document from a collection of reference documents stored in the database. Based on the methodology employed to acquire the data it can be broadly grouped into offline and online [6]. In the former, the temporal attributes such as coordinates, pressure, and angle is recorded by specific devices such as a tablet. As contrast to it, in the latter, the document containing handwritten text is captured in the form of an image and analyzed subsequently. Based on the content of the handwritten text, writer identification systems are categorized into text-dependent and text-independent approaches. In the text-dependent approach, the content of the handwritten text is fixed, while, in the text-independent approach, the content of the handwritten text can vary.
Traditional methods for writer identification rely on extracting handcrafted features to analyze and capture the unique characteristics of a writer’s style from handwritten documents. These features are specifically designed to highlight attributes such as curvature [7], slant [8], and ink width [9], among others. Since these handcrafted features depend on utilizing statistical information from handwriting, they typically require a significant amount of textual data, such as a paragraph or a block of text containing multiple sentences [3, 10] to achieve reliable performance.
Writer identification systems utilizing handcrafted features achieve high recognition rates on public datasets. However, there may be applications for which access to handwritten data is limited. In such cases, decisions may need to be made based on a very small amount of handwriting, such as individual words. This scenario presents a significant challenge, as the writer-specific style information is reduced compared to page or text-level inputs. Consequently, handcrafted feature-based systems fail in these situations, yielding low recognition rates.
@@ II Research framework
Convolutional networks that rely solely on classification-based frameworks typically perform well when distinguishing between dissimilar samples. However, for fine-grained tasks such as writer identification, these approaches often struggle to capture the subtle variations between similar handwriting styles—especially when working with short word images that contain limited textual content. Keeping this in perspective, we introduce a dual stream approach that integrates classification and dissimilarity learning frameworks on the word fragments to achieve more discriminative feature representations. While the classification branch captures localized, writer-specific features from the fragments, the dissimilarity branch focuses on learning pairwise similarities, enabling the model to distinguish fine-grained variations that lie beyond predefined class boundaries. For this, we consider utilizing a Siamese architecture that is first pre-trained on a writer independent dataset but then subsequently fine-tuned using the fragments of the words. An integrated attention mechanism further enhances our framework by highlighting the most informative regions within each fragment. This attention-guided, learning strategy on the fragments significantly improves the ability of the model to identify subtle cues, leading to better performance in writer identification tasks.
@@
1) We suggest a proposal that integrates classification and dissimilarity frameworks in a dual channel network to extract both localized, writer-specific features and writer-independent features (that go beyond predefined class boundaries). More specifically, as we shall see in the sub-section III-A, the writer-dependent module focuses on capturing fine-grained characteristics specific to an individual writer at the fragment level. In contrast, the writer-independent module aims to learn representations of fragments by leveraging the shared characteristics contributed by all writers enrolled in the system.
@@ III Proposed Method
In this section, we present the details of the proposed method for writer identification. Our strategy leverages a framework based on Convolutional Neural Networks and attention mechanism. We process image fragments extracted from word images through two parallel CNN modules: one trained in a writer-dependent manner to capture writer-specific details from the fragments and the other employing a writer-independent approach to extract features of fragments that are prevalent across several writers. The combined features (post attention) are passed through a classification block with fully connected layers, which assign probability scores to the fragments using a softmax activation function. The scores are then accumulated across all fragments of a word image, and the identity is determined based on the maximum accumulated score. Further, the attention mechanism leverages relationships within the convolution layer feature maps, highlighting critical regions in the fragments. By generating attention weights, the network refines its ability to discern relevant features, thus improving performance in writer identification.
@@ III-A Proposed Dual-stream CNN Model
The writer-independent module is designed to extract features from input fragments, that tend to remain consistent across different writing instances of the individuals enrolled in the system. These patterns can be incorporated into writer identification, particularly in unconstrained settings.
@@
The Siamese network is pretrained on the Omniglot dataset [18, 19] , which was chosen for its diverse range of character styles and its suitability for one-shot image recognition tasks  [^2] In order to support the claim on the diversity range of character styles of the Omniglot dataset, we compare its performance in sub-section IV-B with the EMNIST dataset.. Architecturally, each of the CNN pathways of the Siamese writer-independent network mirrors the writer-dependent model (presented in Figure 1(b)), with the addition of global average pooling, dropout, and fully connected layer after the third residual block.
The pre-training of the writer independent block is conducted using the TensorFlow framework, with the network optimized via the Adam optimizer [20] over 100 epochs. The learning rate is initialized at 0.001, and a weight decay factor of 0.5 is applied when the validation loss fails to improve over five consecutive epochs. Upon integration with the WD module, we leverage the idea of transfer learning by fine tuning some of the later layers with the fragments of the word.
@@
$\hat{F}=f_{c}(F_{\textit{WD}},F_{WI})$ (1)
Here $F_{WI}$ and $F_{\textit{WD}}$ represents the writer-dependent and independent features, respectively, while $f_{c}$ denotes the combining function. The combined feature (F^)($\hat${F}) is obtained by applying the combining function $f_{c}$ to the feature maps extracted from the outputs of the third residual block of both branches. In our study, we have explored different methodologies to obtain an efficient combining function, the details of which is discussed in section IV-C.
The combined feature maps are then passed to the classification block, which consists of a global average pooling (GAP) layer, a dropout layer, and a Fully Connected (FC) layer. The detailed architecture of the classification block is illustrated in Figure 1(d).
The network in Figure 1 (a) is trained in an end-to-end framework, where the weights of the writer-dependent module are updated from scratch, while the writer-independent module leverages transfer learning. Specifically, the weights up to the second residual block of this module are kept frozen, and fine-tuning is performed on the subsequent layers to adapt the pre-trained parameters to the fragments of the input word.
@@ III-B Attention module
In this work, we propose an attention mechanism to selectively focus on specific elements within the fragments of a word, thus assigning varying levels of importance to different regions. This adaptability boosts the capability of the model to discern complex patterns and relationships, thus enabling the combined network of Figure 1 (a) to learn more informative features.
@@
Our approach utilizes a MobileViT-inspired multi-head attention mechanism, outlined in [21]. The input feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ is reshaped into a two-dimensional matrix $\mathbf{X}_{U}\in\mathbb{R}^{N\times C}$, where $N=H\cdot W$ denote the spatial dimensions of X{X}. An encoder block $G$ projects $\mathbf{X}_{U}$ into a $d$-dimensional embedding space.
@@
After reshaping the resulting decoder output to $\mathbb{R}^{H\times W\times C}$, it is added to the original input feature map $X$, to fuse the local and global contextual representations. Figure 3 shows the pictorial overview of the attention module.
@@ III-C Implementation details
The fragments extracted from the word image serve as input to the unified model, which integrates both the writer-dependent and writer-independent modules. During training, gradient backpropagation is guided by the label smoothing cross-entropy loss function [22], which converts the one-hot encoded hard labels $y_{i}$ into soft labels $\hat{y}_{i}$ as follows:
@@
Each word is divided into fragments of equal size. With regards to each of the fragments, they are resized to $105\times 105$ while maintaining the aspect ratio and padding (if necessary) with white pixels. These modified fragments are then passed through the dual-stream CNN module to obtain scores that are then accumulated to get the identity of the writer. In our implementation, we use nine fragments .
@@ IV-A Dataset Description
In the following subsections, we present an ablation study on each of the modules in the proposed system. The contribution of individual modules is analyzed through a series of experiments, highlighting their impact on the overall performance.
IV-B Performance evaluation of the writer-dependent (WD) and writer independent (WI) module
@@
In the first experiment, we compare the individual performance of the baseline writer-dependent (WD) module with that of the writer-independent (WI) module. The writer-dependent module is trained end-to-end using fragments extracted from the words of a specific writer. In order to obtain the identity of the user from this module alone, we consider the end to end version of Figure 1 (b) by augmenting the blocks of global average pooling, dropout, and a fully connected layer after the last residual network.
@@
In contrast, the writer-independent module leverages the transfer learning idea wherein the layers of the CNN are first pre-trained using a Siamese architecture with a triplet loss function. It may be worth reminding here that the network being implemented is the end to end version of Figure 1 (b) obtained by augmenting the blocks of global average pooling, dropout, and a fully connected layer after the last residual network. Subsequent to pre-training, writer fragments are passed through one branch of the Siamese network and the parameters after the second residual network are fine tuned for evaluation in a multi-label classifier setting  [^3] In this setting, we consider the output layer size to correspond to the number of writers enrolled in the system. using the label smoothing cross entropy loss. In this ablation study, we consider the performance of the fine-tuned WI modules that have been pre-trained on two separate datasets: EMNIST [27] and Omniglot [18]. The primary objective of using two different datasets for pre-training the writer-independent network is to assess the influence of dataset choice on its performance.
@@
Table II presents the performance of the baseline writer-dependent module, while Table III shows the results of the writer-independent module pre-trained first on the EMNIST and Omniglot datasets and subsequently fine tuned. The evaluation of the module is conducted across different dimensions of the embedded feature representations.
@@
Based on the results presented in Tables II and III, it can be observed that the baseline WD module trained directly on writer fragments outperforms the network trained on writer-independent datasets. Moreover, from Table III following observations can be made:
  1. 1.
    The performance of the WI module initially improves as the size of the embedding layer increases with the best accuracy obtained at 512 dimensions. Furthermore, it can be inferred that higher embedding dimensions (such as 1024 and 2048) introduce redundant information that does not contribute meaningfully to the task, resulting in a drop in performance.
  2. 2.
@@
    The WI module pre-trained on the Omniglot dataset outperforms to the one on the EMNIST. This is primarily owing to the handwriting samples of Omniglot that encompass varied character styles, promoting the learning of diverse writer-independent features. This diversity reduces the risk of the network overfitting to dataset-specific characteristics, thereby enhancing its ability to learn better across different writers with varying styles.
In the light of the above aspects, hereinafter, for the rest of the experiments, we consider the pre-trained parameters of the WI module corresponding to an embedding dimension of 512, as obtained from the Omniglot database.
@@
TABLE III: Average writer identification rate (in %) based on the word image fragment representations obtained from the penultimate layer of the WI module. The pre-training of the module is done on the samples of the EMNIST and Omniglot datasets with varying sizes of the penultimate layer. The best identification rate is marked in bold.
@@ IV-C Impact of the dual stream network without attention
Table IV demonstrates that integrating the writer-independent (WI) module with the baseline writer-dependent (WD) model leads to a notable improvement in overall system performance. This enhancement can be attributed to the complementary strengths of both modules. The writer-dependent component is adept at capturing fine-grained, writer-specific local features by training directly on fragments from the target writers. However, its ability to learn diverse patterns is often limited, and constrained by the amount of text available for training. However upon integration with the writer-independent module, the resulting combined model is able to capture cues that encompass variations in writing style and character morphology.
@@ IV-D Impact of the dual stream network with attention
We now investigate the impact of incorporating an attention module on the overall performance of the system, as discussed in section III-B. Explicitly, two configurations of the attention module presented in Fig 2 (a) and (b) are analyzed using the WI+WD+Concat framework.
Table V summarizes the results obtained from these configurations. Based on the entries, it may be inferred that integrating the attention mechanism impacts the overall performance of the system. Upon comparing the baseline with the attention configuration in Fig 2 (a), the identification precision shows notable improvements with Top 1 accuracy increasing from 92.07 %, 89.76 %, 96.05 % to 93.92 %, 92.62 % and 97.71 %, for the IAM, CVL, and CERUG-EN datasets, respectively. Likewise, the configuration in Fig 2 (b) results in best accuracy of 94.43 %, 93.15 % and 97.86 %. The results pertain to a multi-head attention setup with 2 attention heads and a dimensionality of 64 for the query, key, and value vectors. This configuration is chosen based on empirical observation that increasing the number of heads does not improve overall system performance. Additionally, increasing the dimension of the query vectors tends to reduce performance, suggesting that a compact representation is more effective.
@@
TABLE VI: Performance of writer identification system with WD module trained on word and fragments.
@@
IV-E Efficacy of fragment driven training of the dual network
Moving further, we would like to investigate how the baseline CNN network (WD module) would perform when trained on whole word images in place of its constituent fragments. It is interesting to see that the fragment level training of the network outperforms over that of whole word images (refer Table VI). This trend may be attributed to the fact that convolution networks trained on word images make decisions based on the holistic features, which may not fully capture all the relevant details of the writing styles. Contrast to that, our fragment-based training approach treats each image as a collection of local patches, thereby possibly allowing the network to learn more relevant features. This results in an enhanced representation and in turn explains the effectiveness of the proposed identification process.
For completeness, we also present the influence of CNNs trained on both word images and image fragments using heat map visualizations, corresponding to a sample from the IAM, CVL and CERUG-EN dataset (Figure 5). These heat maps highlight the regions of the input emphasized by the network. Specifically, column 2 shows the activation map from the baseline WD module trained on complete word images, while columns (3–5) display activation maps generated by models trained on image fragments under (i) the WD module, (ii) combined (WD+WI+Concat) module without attention and (iii) combined (WD+WI+Concat) module with the attention configuration of Fig 2(b), respectively. Red regions indicate higher attention focus by the network, while blue regions represent lower attention. Note that all heatmaps are post-processed using Gaussian smoothing to reduce blockiness introduced by fragment-wise reconstruction and to enhance visual coherence.
The model trained on complete word images (column 2) produces focused activations only on selective parts of the word, while some regions get unnoticed. Contrast to it, the attention-based configuration (column 5) yields well focused activations, that closely align with handwriting strokes and highlight discriminative regions in the fragments, thereby demonstrating its ability to capture writer features. Further, the higher attention in the heatmaps in Column 5 aligns with the improved performance of the proposed architecture of Figure 1 (a) , when trained solely by using the fragments of the word.
Figure 5: Visualization of heatmap activation across different network configurations. Rows 1–3 correspond to a word sample from the the IAM, CVL, and CERUG-EN datasets, respectively. Column 2 shows the heatmap produced by the WD module trained on complete word images, while columns 3–5 display heatmaps obtained from networks trained with the (i) the WD module, (ii) combined (WD+WI+Concat) module without attention and (iii) combined (WD+WI+Concat) module with the attention configuration of Fig 2(b), respectively.